if sin(x)=nsin(x+2a) then find the value oftan(x+a)
 

Grade:12th pass

1 Answers

Himanshu
7 years ago
  • sin(x+2a) / sin(x) = 1/n
  • [sin(x+2a) + sin(x)] / [sin(x+2a) – sin(x)] = (1+n) / (1-n)
(:.Using componendo and dividendo)
  • 2.sin(x+a).cos(a) / 2.sin(a).cos(x+a) = (1+n) / (1-n)
  • tan(x+a) / tan(a) = (1+n) / (1-n)
  • tan(x+a) = (1+n) × tan(a)/ (1-n)
The above is the required answer.